How many 5 MeV alpha particles are required to deposit a total energy of 1 J?


1
Expert's answer
2021-06-14T08:19:06-0400

First, remember what 1 MeV is: 1,000,000 eV. What is 1 eV? The energy required to move an electrons of elementary charge e in the potential of 1 V:

1 eV is 1.609·10-19·1=1.609·10-19 J. Therefore:


"N=\frac{1}{5000000·1.609·10^{-19}}=1.25·10^{12}."